If two cards are drawn without replacement from an ordinary​ deck, find the probability of a jack and a 3 being drawn.
The probability of a jack and a 3 being drawn is

Found 2 solutions by addingup, greenestamps:
Answer by addingup(3677)

P(J, then 3) = (1/13)(4/51) = 4/663 = 0.6% probability

Answer by greenestamps(13195)



For the problem exactly as stated, the solution by the other tutor is incorrect.

The problem does not say a jack is drawn first and a 3 is drawn second; it says the two cards drawn are a jack and a 3.

P(J,3) = (4/52)(4/51) = (1/13)(4/51) = 4/663
P(3,J) = (4/52)(4/51) = (1/13)(4/51) = 4/663

P(jack and 3 in two draws) = 8/663